# Estimates of automorphic forms on $\mathrm{SU}(n,1)$

### v1

@@ Abstract.
$\sup_{z\in\overline{X}_{\Gamma}}\bigg{|}\frac{\mu_{\mathrm{Ber},k}^{\mathrm{vol}}(z)}{\mu_{\mathrm{hyp}}^{\mathrm{vol}}(z)}\bigg{|}=O_{X_{\Gamma}}\big{(}k^{4n-1}\big{)},$
@@ 1.2. Statement of results
(2) $\sup_{z\in\overline{X}_{\Gamma}}\bigg{|}\frac{\mu_{\mathrm{Ber},k}^{\mathrm{vol}}(z)}{\operatorname{\mu_{hyp}^{vol}}(z)}\bigg{|}=O_{X_{\Gamma}}\big{(}k^{4n-1}\big{)},$
@@
Our estimate of the Bergman metric (2) is far from being optimal. As per evidence from the one-dimensional setting from [AM23], we conjecture the following estimate
$\sup_{z\in\overline{X}_{\Gamma}}\bigg{|}\frac{\mu_{\mathrm{Ber},k}^{\mathrm{vol}}(z)}{\operatorname{\mu_{hyp}^{vol}}(z)}\bigg{|}=O_{X_{\Gamma}}\big{(}k^{3n}\big{)},$
@@ 2.1. The ball model
The space is endowed with its usual Kähler-Bergman metric $\mu_{\mathrm{hyp}}$, which has constant negative curvature $-1$. For any given $z=(z_{1},\ldots,z_{n})\in\mathbb{B}^{n}$, the hyperbolic metric $\mu_{\mathrm{hyp}}(z)$ is given by the following formula
@@ 2.2. Alternate models of $\mathcal{H}^{n}(\mathbb{C})$
The definition of $z\in\mathbb{D}^{n}$ can be rephrased as $-2\mathrm{Im}(z_{n})<-\sum_{j=1}^{n-1}|z_{j}|^{2}$. This ensures that $\gamma_{21}(z)\in\mathbb{B}^{n}$.
@@
(9) $(z_{1},\dots,z_{n})$ $\mapsto\left(\frac{z_{1}+1}{z_{1}-1},\frac{z_{2}}{z_{1}-1},\dots,\frac{z_{n}}{z_{1}-1}\right).$
It is again a straightforward check that for $z\in\mathbb{E}^{n}$, we have $\gamma_{31}(z)\in\mathbb{B}^{n}$.
Finally we define $\gamma_{23}=\gamma_{13}\circ\gamma_{21}$, where $\gamma_{13}=\gamma^{-1}_{31}$, and $\gamma_{23}$ is an isometry $(\mathbb{D}^{n},H_{2})\to(\mathbb{E}^{n},H_{3})$ between models (2) and (3).
@@ 2.3. Picard variety
Here, $z$ is seen as a column vector vector in $\mathbb{C}^{n}$.
@@
Locally, we identify the Picard variety $X_{\Gamma}$ with its universal cover $\mathbb{B}^{n}$. Locally, for $z,w\in X_{\Gamma}$ (identifying $X_{\Gamma}$ with $\mathbb{B}^{n}$), the hyperbolic distance between the points $z$ and $w$, is given by formula (7). Furthermore, let $\infty=(1,0,\ldots,0)^{t}\in\mathcal{F}_{\Gamma}$ denote a fixed fundamental domain of $X_{\Gamma}$, and let ¯$\overline{\mathcal{F}}_{\Gamma}:=\mathcal{F}_{\Gamma}\cup\{\infty\}$.
@@ 2.4. Stabilizers of the cusps and cuspidal neighborhood of $X_{\Gamma}$
[…]
@@ 2.5. Determinant bundle of the Cotangent bundle
Let $\overline{X}_{\Gamma}:=X_{\Gamma}\cup\{\infty\}$ denote the one-point compactification of $X_{\Gamma}$. Let $\Lambda:=\Omega^{n}_{\overline{X}_{\Gamma}}$ denote the line bundle of holomorphic $(n,0)$-forms on $\overline{X}_{\Gamma}$. Furthermore, let $\lambda^{k}:=\Lambda^{\otimes k}\otimes O_{\overline{X}_{\Gamma}}((k-1)\infty)$, denote the space of meromorphic differential forms, with a pole of order $k-1$ at the cusp $\infty$, and which remain holomorphic on $\overline{X}_{\Gamma}\backslash\{\infty\}$.
@@ 3.1. Estimates of the Bergman kernel
Applying estimates (27) and (2.8), and adapting arguments from Theorem 4.2 in [ABR23] to the compact subset $\mathcal{A}\subset\mathcal{F}_{\Gamma}$, we arrive at the following estimate
$\sum_{\gamma\in\Gamma}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}}{\cosh^{k(n+1)}(\operatorname{d_{\mathrm{hyp}}}(z,\gamma z)/2)}\leq$
(33) $\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}+\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\cosh^{2n}(r_{\mathcal{A}}/4)}{(k(n+1)-2n-1)\sinh^{2n}(r_{\mathcal{A}}/4)}+\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\sin^{2n}(5r_{\mathcal{A}}/8)}{\sinh^{2n}(r_{\mathcal{A}}/4)\cosh^{k(n+1)}(3r_{\mathcal{A}}/8)},$
(34) $\mathrm{where}\,\,r_{\mathcal{A}}:=\big{\{}\operatorname{d_{\mathrm{hyp}}}(z,\gamma z)\big{|}\,z\in\mathcal{A},\,\,\gamma\in\Gamma\backslash\{\mathrm{Id}\}\big{\}},$
@@
(35) $\big{|}\operatorname{{\it{\mathcal{B}}}_{{\it{X}}_{\Gamma}}^{\lambda^{{\it k}}}}(z,z)\big{|}_{\mathrm{hyp}}\geq\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}-\sum_{\gamma\in\Gamma\backslash\{\mathrm{Id}\}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}}{\cosh^{k(n+1)}(\operatorname{d_{\mathrm{hyp}}}(z,\gamma z)/2)},$
@@
From arguments used tp prove Theorem 4.2 in [ABR23], we have the following estimate
$\sum_{\gamma\in\Gamma\backslash\{\mathrm{Id}\}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}}{\cosh^{k(n+1)}(\operatorname{d_{\mathrm{hyp}}}(z,\gamma z)/2)}\leq$
(36) $\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\cosh^{2n}(r_{\mathcal{A}}/4)}{(k(n+1)-2n-1)\sinh^{2n}(r_{\mathcal{A}}/4)}+\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\sin^{2n}(5r_{\mathcal{A}}/8)}{\sinh^{2n}(r_{\mathcal{A}}/4)\cosh^{k(n+1)}(3r_{\mathcal{A}}/8)},$
@@
$\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}-\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\cosh^{2n}(r_{\mathcal{A}}/4)}{(k(n+1)-2n-1)\sinh^{2n}(r_{\mathcal{A}}/4)}-\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\sin^{2n}(5r_{\mathcal{A}}/8)}{\sinh^{2n}(r_{\mathcal{A}}/4)\cosh^{k(n+1)}(3r_{\mathcal{A}}/8)}\leq$
$\big{|}\operatorname{{\it{\mathcal{B}}}_{{\it{X}}_{\Gamma}}^{\lambda^{{\it k}}}}(z,z)\big{|}_{\mathrm{hyp}}=O_{\mathcal{A}}\big{(}k^{n}\big{)},$
@@
[…]
@@
Observing that the function $P(z)$ approaches zero, as $x_{1}$ approaches $\infty$- ∞, we infer that it attains its maxima on the set $\partial U_{4\pi/k(n+1)}(\infty)$, which together with equation (38), completes the proof of the lemma. ∎
@@
$\sup_{z\in\mathcal{A}_{\epsilon_{0}}}\big{|}\operatorname{{\it{\mathcal{B}}}_{{\it{X}}_{\Gamma}}^{\lambda^{{\it k}}}}(\gamma_{31}z,\gamma_{31}z)\big{|}_{\mathrm{hyp}}=\sup_{z\in\mathcal{A}_{\epsilon_{0}}}\sum_{\gamma\in\Gamma_{3}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}}{\cosh^{k(n+1)}(\operatorname{d_{\mathrm{hyp}}}(z,\gamma z)/2)}\leq$
(40) $\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\cosh^{2n}(r_{\epsilon_{0}}/4)}{(k(n+1)-2n-1)\sinh^{2n}(r_{\epsilon_{0}}/4)}+\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\sin^{2n}(5r_{\epsilon_{0}}/8)}{\sinh^{2n}(r_{\epsilon_{0}}/4)\cosh^{k(n+1)}(3r_{\epsilon_{0}}/8)}=O_{X_{\Gamma}}\big{(}k^{n}\big{)},$
$\mathrm{where}\,\,r_{\epsilon_{0}}:=\inf\big{\{}\operatorname{d_{\mathrm{hyp}}}(z,\gamma z)\big{|}\,z\in\mathcal{A}_{\epsilon_{0}},\,\,\gamma\in\Gamma_{3}\backslash\{\mathrm{Id}\}\big{\}}.$
@@
(43) $\langle\tilde{\gamma z},\tilde{z}\rangle_{H_{3}}=-\frac{k(n+1)}{4\pi}+\frac{-|\tau|^{2}+it}{2};\,\,\langle\tilde{z},\tilde{\gamma z}\rangle_{H_{3}}=-\frac{k(n+1)}{4\pi}+\frac{-|\tau|^{2}-it}{2},$
@@
(45) $\sum_{u_{n}\in\mathbb{Z}}\ldots\sum_{u_{1}\in\mathbb{Z}}\sum_{v_{n}\in\mathbb{Z}}\ldots\sum_{v_{1}\in\mathbb{Z}}\sum_{t\in\mathbb{Z}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}\big{(}k(n+1)/4\pi+\sum_{j=1}^{n}(u_{j}^{2}+v_{j}^{2})\big{)}^{2}+t^{2}\big{)}^{k(n+1)/2}}.$
@@
$\sum_{t\in\mathbb{Z}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}\big{(}k(n+1)/4\pi+\sum_{j=1}^{n}(u_{j}^{2}+v_{j}^{2})\big{)}^{2}+t^{2}\big{)}^{k(n+1)/2}}\leq$
$\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}k(n+1)/4\pi+\sum_{j=1}^{n}(u_{j}^{2}+v_{j}^{2})\big{)}^{k(n+1)}}+$
(46) $\int_{0}^{\infty}\frac{2\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}dt}{\big{(}\big{(}k(n+1)/4\pi+\sum_{j=1}^{n}(u_{j}^{2}+v_{j}^{2})\big{)}^{2}+t^{2}\big{)}^{k(n+1)/2}}.$
@@
$\int_{0}^{\infty}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}dt}{\big{(}\big{(}k(n+1)/4\pi+\sum_{j=1}^{n}(u_{j}^{2}+v_{j}^{2})\big{)}^{2}+t^{2}\big{)}^{k(n+1)/2}}=$
$\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}k(n+1)/4\pi+\sum_{j=1}^{n}(u_{j}^{2}+v_{j}^{2})\big{)}^{k(n+1)-1}}\int_{0}^{\infty}\frac{d\theta}{\big{(}\theta^{2}+1\big{)}^{k(n+1)/2}}\ll$
(49) $\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}k(n+1)/4\pi+\sum_{j=1}^{n}(u_{j}^{2}+v_{j}^{2})\big{)}^{k(n+1)-1}}\cdot\frac{\Gamma((k(n+1)-1)/2)}{\Gamma(k(n+1)/2)}.$
@@
$\sum_{u_{n}\in\mathbb{Z}}\ldots\sum_{u_{1}\in\mathbb{Z}}\sum_{v_{n}\in\mathbb{Z}}\ldots\sum_{v_{1}\in\mathbb{Z}}\sum_{t\in\mathbb{Z}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}\big{(}k(n+1)/4\pi+\sum_{j=1}^{n}(u_{j}^{2}+v_{j}^{2})\big{)}^{2}+t^{2}\big{)}^{k(n+1)/2}}\ll$
$\sum_{u_{n}\in\mathbb{Z}}\ldots\sum_{u_{1}\in\mathbb{Z}}\sum_{v_{n}\in\mathbb{Z}}\ldots\sum_{v_{1}\in\mathbb{Z}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}k(n+1)/4\pi+\sum_{j=1}^{n}(u_{j}^{2}+v_{j}^{2})\big{)}^{k(n+1)}}+$
(50) $\sum_{u_{n}\in\mathbb{Z}}\ldots\sum_{u_{1}\in\mathbb{Z}}\sum_{v_{n}\in\mathbb{Z}}\ldots\sum_{v_{1}\in\mathbb{Z}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}k(n+1)/4\pi+\sum_{j=1}^{n}(u_{j}^{2}+v_{j}^{2})\big{)}^{k(n+1)-1}}\cdot\frac{\Gamma((k(n+1)-1)/2)}{\Gamma(k(n+1)/2)}.$

### v2

@@ Abstract.
$\sup_{z\in\overline{X}_{\Gamma}}\bigg{|}\frac{\mu_{\mathrm{Ber},k}^{\mathrm{vol}}(z)}{\mu_{\mathrm{hyp}}^{\mathrm{vol}}(z)}\bigg{|}=O_{X_{\Gamma}}\big{(}k^{2(n-1)(n+2)+n+3}\big{)},$
@@ 1.1. History
Using a different set of techniques, a similar estimate as our first main result was derived by Zhou in [Zh24]. The approach in [Zh24] is an application of the method of peak sections by Tian.
@@ 1.2. Statement of results
(2) $\sup_{z\in\overline{X}_{\Gamma}}\bigg{|}\frac{\mu_{\mathrm{Ber},k}^{\mathrm{vol}}(z)}{\operatorname{\mu_{hyp}^{vol}}(z)}\bigg{|}=O_{X_{\Gamma}}\big{(}k^{2(n-1)(n+2)+n+3}\big{)},$
@@
Remark 1.2.
Our estimate of the Bergman metric (2) is far from being optimal. The power of $k$ in our estimate (2) is quadratic in $n$, and we believe it should linear in $n$. For $n\geq 1$, we conjecture the following estimate
$\sup_{z\in\overline{X}_{\Gamma}}\bigg{|}\frac{\mu_{\mathrm{Ber},k}^{\mathrm{vol}}(z)}{\operatorname{\mu_{hyp}^{vol}}(z)}\bigg{|}=O_{X_{\Gamma}}\big{(}k^{4n}\big{)},$
@@
As per evidence from the one-dimensional setting from [AMW25], our estimate (2) is optimal for the case $n=1$. It is to be mentioned that the lower bound proved in Proposition 3.6, corrects and completes arguments proved in [AM23], for the case $n=1$. Hence, for $n=1$, we have the following estimate
$\sup_{z\in\overline{X}_{\Gamma}}\bigg{|}\frac{\mu_{\mathrm{Ber},k}^{\mathrm{vol}}(z)}{\operatorname{\mu_{hyp}^{vol}}(z)}\bigg{|}=O_{X_{\Gamma}}\big{(}k^{4}\big{)},$
@@ 2.1. The ball model
The space is endowed with its usual Kähler-Bergman metric $\mu_{\mathrm{hyp}}$, which has constant negative curvature 1-1. For any given $z=(z_{1},\ldots,z_{n})\in\mathbb{B}^{n}$, the hyperbolic metric $\mu_{\mathrm{hyp}}(z)$ is given by the following formula
@@ 2.2. Alternate models of $\mathcal{H}^{n}(\mathbb{C})$
The definition of $z\in\mathbb{D}^{n}$ can be rephrased as $2\mathrm{Im}(z_{n})<-\sum_{j=1}^{n-1}|z_{j}|^{2}$. This ensures that $\gamma_{21}(z)\in\mathbb{B}^{n}$.
@@
(9) $(z_{1},\dots,z_{n})$ $\mapsto\left(\frac{z_{1}+1}{z_{1}-1},\frac{\sqrt{2}z_{2}}{z_{1}-1},\dots,\frac{\sqrt{2}z_{n}}{z_{1}-1}\right).$
It is again a straightforward check that for $z\in\mathbb{E}^{n}$, we have $\gamma_{31}(z)\in\mathbb{B}^{n}$, and the inverse map is given by
$\gamma_{13}:(\mathbb{B}^{n},H_{1})$ $\longrightarrow(\mathbb{E}^{n},H_{3})$
(10) $(z_{1},\dots,z_{n})$ $\mapsto\left(\frac{z_{1}+1}{z_{1}-1},\frac{\sqrt{2}z_{2}}{z_{1}-1},\dots,\frac{\sqrt{2}z_{n}}{z_{1}-1}\right).$
Finally we define $\gamma_{23}=\gamma_{13}\circ\gamma_{21}:(\mathbb{D}^{n},H_{2})\to(\mathbb{E}^{n},H_{3})$, which is an isometry between models (2) and (3).
@@ 2.3. Picard variety
Here, $z$ is seen as a column vector in $\mathbb{C}^{n}$.
@@
Locally, we identify the Picard variety $X_{\Gamma}$ with its universal cover $\mathbb{B}^{n}$. Locally, for $z,w\in X_{\Gamma}$ (identifying $X_{\Gamma}$ with $\mathbb{B}^{n}$), the hyperbolic distance between the points $z$ and $w$, is given by formula (7). Furthermore, let $\infty=(1,0,\ldots,0)^{t}\in\mathcal{F}_{\Gamma}$ denote a fixed fundamental domain of $X_{\Gamma}$, and let $\overline{\mathcal{F}}_{\Gamma}:=\mathcal{F}_{\Gamma}\cup\{\infty\}$.
@@ 2.4. Stabilizers of the cusps and cuspidal neighborhood of $X_{\Gamma}$
Remark 2.1.
Let us make the group $\Gamma_{3,\infty}$ precise when $\Gamma_{3}=\mathrm{SU}((n,1),\mathcal{O}_{K})$. We have $\tau\in\mathcal{O}_{K}^{n-1}$ and $U\in\mathrm{SU}(n-1,\mathcal{O}_{K})$ as before. We only need to determine a more precise condition on $t$. Since $\tau\in\mathcal{O}_{K}^{n-1}$, we have $|\tau|^{2}\in\mathcal{O}_{K}\cap\mathbb{R}_{+}=\mathbb{Z}_{+}$. Therefore, the condition $\frac{-|\tau|^{2}+it}{2}\in\mathcal{O}_{K}$ is equivalent to $it\in\mathcal{O}_{K}$ and $|\tau|^{4}+t^{2}\in 4\mathbb{Z}$.
Now, if we take $K=\mathbb{Q}[\sqrt{-d}]$, then we have $\mathcal{O}_{K}=\mathbb{Z}[\sqrt{-d}]$ if $d\equiv 1\textrm{ or }2\pmod{4}$ and $\mathcal{O}_{K}=\mathbb{Z}[\frac{1+\sqrt{-d}}{2}]$ if $d\equiv 3\pmod{4}$. In either case, the condition $it\in\mathcal{O}_{K}$ is equivalent to $t=b\sqrt{d}$ for some $b\in\mathbb{Z}$. When $d\equiv 3\pmod{4}$, the second condition implies that the set of all possible $\tau$ is full lattice $\mathcal{O}_{K}^{n-1}$ and all possible $t$ is $\mathbb{Z}\sqrt{d}$. When $d\equiv 1\textrm{ or }2\pmod{4}$, the second condition implies that the set of all possible $\tau$ is the sublattice of index $2$ in $\mathcal{O}_{K}^{n-1}$ given by $\{\tau\in\mathcal{O}_{K}^{n-1}\mid|\tau|^{2}\in 2\mathbb{Z}\}$ and all possible $t$ is $2\mathbb{Z}\sqrt{d}$.
@@
$U_{\epsilon}(\infty):=\big{\{}z:=(z_{1},\ldots,z_{n})^{t}\in\mathcal{F}_{\Gamma_{3}}\big{|}\,2\mathrm{Re}(z_{1})\leq-1/\epsilon,\,\,z_{2}=\cdots=z_{n}=0\big{\}};$
@@ 2.5. Determinant bundle of the Cotangent bundle
[…]
@@ 3.1. Estimates of the Bergman kernel
Applying and adatping estimates (28) and (2.8) (refining estimate (50) in [ABR23]), and adapting arguments from Theorem 4.2 in [ABR23] to the compact subset $\mathcal{A}\subset\mathcal{F}_{\Gamma}$, we arrive at the following estimate
$\sum_{\gamma\in\Gamma}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}}{\cosh^{k(n+1)}(\operatorname{d_{\mathrm{hyp}}}(z,\gamma z)/2)}\leq\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}+\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\sin^{2n}(5r_{\mathcal{A}}/8)}{\sinh^{2n}(r_{\mathcal{A}}/4)\cosh^{k(n+1)}(3r_{\mathcal{A}}/8)}+$
(34) $\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\cosh^{2n}(r_{\mathcal{A}}/4)}{(k(n+1)-2n-1)\sinh^{2n}(r_{\mathcal{A}}/4)\cosh^{k(n+1)-2n-2}(3r_{\mathcal{A}}/8)},$
[…]
@@
[…]
@@
From arguments as in Proposition 3.1 (estimate (3.1)), we have the following estimate
[…]
(37) $\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\cosh^{2n}(r_{\mathcal{A}}/4)}{(k(n+1)-2n-1)\sinh^{2n}(r_{\mathcal{A}}/4)\cosh^{k(n+1)-2n-2}(3r_{\mathcal{A}}/8)},$
@@
$\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}-\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\sin^{2n}(5r_{\mathcal{A}}/8)}{\sinh^{2n}(r_{\mathcal{A}}/4)\cosh^{k(n+1)}(3r_{\mathcal{A}}/8)}$-
$\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\cosh^{2n}(r_{\mathcal{A}}/4)}{(k(n+1)-2n-1)\sinh^{2n}(r_{\mathcal{A}}/4)\cosh^{k(n+1)-2n-2}(3r_{\mathcal{A}}/8)}\leq\big{|}\operatorname{{\it{\mathcal{B}}}_{{\it{X}}_{\Gamma}}^{\lambda^{{\it k}}}}(z,z)\big{|}_{\mathrm{hyp}}=O_{\mathcal{A}}\big{(}k^{n}\big{)},$
@@
$\frac{\partial P(z)}{\partial x_{i}}=-2k(n+1)x_{i}e^{4\pi x_{1}}\big{(}-2x_{1}-\sum_{i=2}^{n}x_{i}^{2}-\sum_{i=2}^{n}y_{i}^{2}\big{)}^{k(n+1)-1},\,\,\mathrm{for\,\,all}\,\,2\leq i\leq n;$
$\frac{\partial P(z)}{\partial y_{i}}=-2k(n+1)y_{i}e^{4\pi x_{1}}\big{(}-2x_{1}-\sum_{i=2}^{n}x_{i}^{2}-\sum_{i=2}^{n}y_{i}^{2}\big{)}^{k(n+1)-1},\,\,\mathrm{for\,\,all}\,\,1\leq i\leq n.$
@@
Observing that the function $P(z)$ approaches zero, as $x_{1}$ approaches -$\infty$, we infer that it attains its maxima on the set $\partial U_{4\pi/k(n+1)}(\infty)$, which together with equation (39), completes the proof of the lemma. ∎
@@
$\sup_{z\in\mathcal{A}_{\epsilon_{0}}}\big{|}\operatorname{{\it{\mathcal{B}}}_{{\it{X}}_{\Gamma}}^{\lambda^{{\it k}}}}(\gamma_{31}z,\gamma_{31}z)\big{|}_{\mathrm{hyp}}\leq\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}+\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\sin^{2n}(5r_{\epsilon_{0}}/8)}{\sinh^{2n}(r_{\epsilon_{0}}/4)\cosh^{k(n+1)}(3r_{\epsilon_{0}}/8)}+$
$\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\cosh^{2n}(r_{\epsilon_{0}}/4)}{(k(n+1)-2n-1)\sinh^{2n}(r_{\epsilon_{0}}/4)\cosh^{k(n+1)-2n-2}(3r_{\epsilon_{0}}/8)}=O_{X_{\Gamma}}\big{(}k^{n}\big{)},$
[…]
@@
$\langle\tilde{\gamma z},\tilde{z}\rangle_{H_{3}}=2\mathrm{Re}(z_{1})+\sum_{j=2}^{n}|z_{2}|^{2}+\frac{-|\tau|^{2}+it}{2}=-\frac{k(n+1)}{4\pi}+\frac{-|\tau|^{2}+it}{2};$
(44) $\langle\tilde{z},\tilde{\gamma z}\rangle_{H_{3}}=2\mathrm{Re}(z_{1})+\sum_{j=2}^{n}|z_{2}|^{2}+\frac{-|\tau|^{2}-it}{2}-\frac{k(n+1)}{4\pi}+\frac{-|\tau|^{2}-it}{2},$
@@
(46) $\sum_{u_{n-1}\in\mathbb{Z}}\ldots\sum_{u_{1}\in\mathbb{Z}}\sum_{v_{n-1}\in\mathbb{Z}}\ldots\sum_{v_{1}\in\mathbb{Z}}\sum_{t\in\mathbb{Z}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}\big{(}k(n+1)/4\pi+\sum_{j=1}^{n-1}(u_{j}^{2}+v_{j}^{2})\big{)}^{2}+t^{2}\big{)}^{k(n+1)/2}}.$
@@
$\sum_{t\in\mathbb{Z}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}\big{(}k(n+1)/4\pi+\sum_{j=1}^{n-1}(u_{j}^{2}+v_{j}^{2})\big{)}^{2}+t^{2}\big{)}^{k(n+1)/2}}\leq$
$\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}k(n+1)/4\pi+\sum_{j=1}^{n-1}(u_{j}^{2}+v_{j}^{2})\big{)}^{k(n+1)}}+$
(47) $\int_{0}^{\infty}\frac{2\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}dt}{\big{(}\big{(}k(n+1)/4\pi+\sum_{j=1}^{n-1}(u_{j}^{2}+v_{j}^{2})\big{)}^{2}+t^{2}\big{)}^{k(n+1)/2}}.$
@@
$\int_{0}^{\infty}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}dt}{\big{(}\big{(}k(n+1)/4\pi+\sum_{j=1}^{n-1}(u_{j}^{2}+v_{j}^{2})\big{)}^{2}+t^{2}\big{)}^{k(n+1)/2}}=$
$\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}k(n+1)/4\pi+\sum_{j=1}^{n-1}(u_{j}^{2}+v_{j}^{2})\big{)}^{k(n+1)-1}}\int_{0}^{\infty}\frac{d\theta}{\big{(}\theta^{2}+1\big{)}^{k(n+1)/2}}\ll$
(50) $\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}k(n+1)/4\pi+\sum_{j=1}^{n-1}(u_{j}^{2}+v_{j}^{2})\big{)}^{k(n+1)-1}}\cdot\frac{\Gamma((k(n+1)-1)/2)}{\Gamma(k(n+1)/2)}.$
@@
$\sum_{u_{n-1}\in\mathbb{Z}}\ldots\sum_{u_{1}\in\mathbb{Z}}\sum_{v_{n-1}\in\mathbb{Z}}\ldots\sum_{v_{1}\in\mathbb{Z}}\sum_{t\in\mathbb{Z}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}\big{(}k(n+1)/4\pi+\sum_{j=1}^{n-1}(u_{j}^{2}+v_{j}^{2})\big{)}^{2}+t^{2}\big{)}^{k(n+1)/2}}\ll$
$\sum_{u_{n-1}\in\mathbb{Z}}\ldots\sum_{u_{1}\in\mathbb{Z}}\sum_{v_{n-1}\in\mathbb{Z}}\ldots\sum_{v_{1}\in\mathbb{Z}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}k(n+1)/4\pi+\sum_{j=1}^{n-1}(u_{j}^{2}+v_{j}^{2})\big{)}^{k(n+1)}}+$
(51) $\sum_{u_{n-1}\in\mathbb{Z}}\ldots\sum_{u_{1}\in\mathbb{Z}}\sum_{v_{n-1}\in\mathbb{Z}}\ldots\sum_{v_{1}\in\mathbb{Z}}\frac{\operatorname{{\it{C_{\mathrm{X_{\Gamma}}}^{k(n+1)}}}}\big{(}k(n+1)/4\pi\big{)}^{k(n+1)}}{\big{(}k(n+1)/4\pi+\sum_{j=1}^{n-1}(u_{j}^{2}+v_{j}^{2})\big{)}^{k(n+1)-1}}\cdot\frac{\Gamma((k(n+1)-1)/2)}{\Gamma(k(n+1)/2)}.$